# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v2

@@ IV-D2 Ablation Study on the Loss Function
In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets. First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregated the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation. The resulting probabilities are then used for target detection. In Fig. 7, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs. It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets. In Table X, we show the detection performance of the reference model. It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets. In particular, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
@@
As summarized in Table XI, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For example, RadarLLM (PA) achieves an average gain of 9.9% over RadarLLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart. In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
TABLE X: Detection performance of reference model on three low SCR datasets.
Reference model $H$ HV VH $V$
Data1 46.68 45.21 46.75 18.48
Data2 34.87 50.75 48.41 42.51
Data3 47.39 59.82 61.85 44.72
TABLE XI: Results of ablation study on the loss function.
Data1 Data2 Data3 AVG
RadarLLM (PA) 52.06(+16.8) 54.87(+5.4) 65.81(+7.5) 57.58(+9.9)
RadarLLM (WCE) 43.51(+8.8) 53.95(+4.9) 64.55(+6.8) 54.84(+7.7)
RadarLLM (CE) 34.73 48.96 57.80 47.16
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
TABLE XII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of patch size.)
Average DR Average IT(s)
RadarLLM (with a patch size of 32) 79.44 5.9259
RadarLLM (with a patch size of 48) 81.96 5.6561
RadarLLM (with a patch size of 64) 79.09 5.6222
@@
We first investigate the impact of patch size on RadarLLM’s performance. The results of the ablation study are summarized in Table XII. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
@@
We then investigated the impact of the number of GPT2 layers on RadarLLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table XIII. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the cross-modal transfer capabilities of the language model, while excess layers suffer from overfitting due to the task-irrelevant network parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
TABLE XIII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers. (RadarLLM (n) represents RadarLLM with n LLM layers.))
Average DR Average IT(s) Total/Training NP(M)
RadarLLM (0) 77.41 5.3484 41.0410/2.4445
RadarLLM (2) 81.29 5.4812 55.2175/2.4506
RadarLLM (4) 81.96 5.6561 69.3932/2.4567
RadarLLM (6) 80.42 5.7331 83.5690/2.4628
RadarLLM (8) 80.01 5.8655 97.7447/2.4690
IV-F Case Study of the Effectiveness of Feature Token Importance
To further demonstrate the effectiveness of the proposed preference-aware loss function, we conduct a case study analyzing the training weight dynamics of feature tokens for two representive sequence features: Short-Time Fourier Transform Magnitude Spectrum (SMS) and Instantaneous Phase (IP) features.
The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, mainly characterized by the feature tokens 26 and 27. As illustrated in Fig. 7, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns. However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs. This adaptive behavior reflects the model’s ability to dynamically reallocate attention from well-learned patterns to more complex or informative features, thus mitigating the risk of overfitting to easily recognizable signal features.
In contrast, the IP feature demonstrates a more complex structure, encompassing both moderately varying patterns associated with target echos (corresponding to tokens 0, 3, and 4) and rapidly fluctuating patterns caused by clutter-induced noise (corresponding to tokens 8, 9, and 10). These six tokens, arranged from left to right, align with the gray-lighted segments in the IP feature map illustrated in Fig. 8. During training, tokens 0, 3, and 4 consistently receive elevated attention weights, typically exceeding 1, indicating the model’s sustained emphasis on these informative yet challenging-to-learn feature tokens. In contrast, tokens 8, 9, and 10 are assigned substantially lower weights, which highlights the model’s ability to suppress excessive learning from noisy and less informative feature tokens.
(1). SMS feature
(2). SMS feature token weight dynamics
(3). IP feature
(4). IP feature token weight dynamics
Figure 8: Dynamics of feature token weights during training: (a) SMS feature token weight dynamics, (b) IP feature token weight dynamics. The graphs show how the attention weights assigned to individual feature tokens evolve over epochs, emphasizing the model’s ability to adaptively focus on informative features and suppress noisy tokens.
V Conclusion and Future Work
In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection. Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their online evaluated learning values in a lightweight form, effectively reducing model overfitting issues in low SCR environments. Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios. These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection. Looking ahead, several promising research directions emerge:
  1. 1.
    Future studies should investigate the use of more sophisticated open source pre-trained LLMs, such as LLaMA2 or LLaMA3 [44]. These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.
  2. 2.
    Although the present approach demonstrates the effectiveness of the selected reference model, there remains considerable potential to explore alternative architectures. Future investigations could examine models such as Mamba [45] or leverage iterative training frameworks, which may further enhance the performance and training efficiency of the reference model.
  3. 3.
    To address the storage demands of using large language models (LLMs), we suggest employing model compression strategies, including quantization, pruning, and knowledge distillation. These methods aim to reduce model size while maintaining performance, facilitating the deployment of LLMs in resource-constrained environments.

### v3

@@ IV-D3 Loss Function
In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets. First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregated the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation. The resulting probabilities are then used for target detection. In Fig. 7, we present the detection rates on test samples from different low-SCR datasets under HH polarization across multiple training epochs. It can be observed that preference-aware loss achieves an improvement of around 1%-12% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets. We also include the performance of the reference model (gray line) for comparison. it can be seen that the optimal detection performance of the fine-tuned PLM-based model outperforms the reference model by approximately 3%-18% across various datasets. Notably, despite this substantial performance gap, employing the reference model to construct the preference-aware loss still leads to significant performance gains, demonstrating that our method effectively realizes a compelling weak-to-strong generalization capability.
@@
As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For example, RadarPLM (PA) achieves an average gain of 9.9% over RadarPLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart. In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
TABLE IX: Results of ablation study on the loss function.
Method Data1 Data2 Data3 AVG
RadarPLM (PA) 52.06 (+16.8) 54.87 (+5.4) 65.81 (+7.5) 57.58 (+9.9)
RadarPLM (WCE) 43.51 (+8.8) 53.95 (+4.9) 64.55 (+6.8) 54.84 (+7.7)
RadarPLM (CE) 34.73 48.96 57.80 47.16
@@
TABLE X: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers. The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions).
RadarPLM Performance under Different Configurations
Patch Size Configuration Average DR (%) Average IT (s) Total / Training NP (M)
32 79.44 5.9259 69.39 / 2.46
48 81.96 5.6561 69.39 / 2.46
64 79.09 5.6222 69.39 / 2.46
GPT2 Layers 0 Layer 77.41 5.3484 41.04 / 2.44
2 Layers 81.29 5.4812 55.22 / 2.45
4 Layers 81.96 5.6561 69.39 / 2.46
6 Layers 80.42 5.7331 83.57 / 2.46
8 Layers 80.01 5.8655 97.74 / 2.47
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
We first investigate the impact of patch size on RadarPLM’s performance. The results of the ablation study are summarized in Table X. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarPLM.
@@
We then investigated the impact of the number of GPT2 layers on RadarPLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table X. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the parameter transfer capabilities of the PLM, while excess layers suffer from overfitting due to the task-irrelevant parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
V Conclusion
In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection. We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization. Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios. These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing. Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.